OpenStudy (anonymous):

Find the Point at which the given lines intersect? A) L1 is a line going through (1,1,0) and parallel to i-j+2k L2 is a line going through (2,0,2) and (1,1,2). B) Find an equation of the plane containing these lines C) Find the equation of the plane equidistant fromt he points (1,1,0) and (2,0,2) Im completely lost this really makes me want to give up on math for good.

OpenStudy (aravindg):

First find the equation of the lines.

OpenStudy (anonymous):

how do i go about doing that?

OpenStudy (anonymous):

All I know is there will be a cross product somwhere in this equation and im not even 100% sure how to cross product

OpenStudy (anonymous):

ok i'll give you a general view..i'll use general variables to give you a conception, you can replace them with the values specific to your query... so Consider an line L and a point P(x0,y0,z0) on L. Direction of this line is determined by a vector v that is parallel to Line L...Do you agree?

OpenStudy (anonymous):

Yea Given in line L1

OpenStudy (anonymous):

Point (1.1.0) is parallel vector (1.-1,2)

OpenStudy (anonymous):

Let P(x,y,z) be any point on the Line Let r0 is the Position vector of point P0 r is the Position vector of point P... Then vector equation of line is given by r=r0+vt Where t is a scalar Let v=<a,b,c> r=<x0,y0, z0>... plugin your values to get an equation for L1

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

Im totally drawing a blank here.

OpenStudy (anonymous):

for L2..(x-x1)/(x2-x1) = (y-y1)/(y2-y1) = (z-z1)/(z2-z1) gives equation of line passing through (x1,y1,z1) and (x2,y2,z2)

OpenStudy (anonymous):

oh?seems you need to know basics...see the attachment

OpenStudy (anonymous):

thanks for this power point Im heading to clas snow but ill look it over and get back to you, thanks for teh help thus far.

OpenStudy (anonymous):

good luck